OpenStudy (nuccioreggie):

3|x-1|=12 Step by step

rebeccaxhawaii (rebeccaxhawaii):

|| make things positive

OpenStudy (oliviagggg):

http://www.cymath.com/answer.php?q=3%7Cx-1%7C%3D12

OpenStudy (alex6799):

\[3\left| x-1 \right|=12\] \[\frac{ 3\left| x-1 \right| }{ 3 }=\frac{ 12 }{ 3 }\] \[\left| x-1 \right|=\frac{ 2^2*3 }{ 3 }\] \[\left| x-1 \right|=2^2\] \[x-1=4 ; -(x-1)=4\] \[x=5; -x+1=4\] \[x=5;-x=3\] \[x=5;x=-3\] \[x \in \left\{ 5,-3 \right\}\]